OpenStudy (anonymous):

how do you calculate the speed of the center of mass of three objects moving away from one another

OpenStudy (unklerhaukus):

vectors !

OpenStudy (anonymous):

\[Vcm = (m1*v1 + m2*v2+....) / (m1+m2+...)\] all the v are vectors here

OpenStudy (anonymous):

so its basicaly the resultant of the three?

OpenStudy (anonymous):



OpenStudy (anonymous):

first - intuition - vcm is in the same direction as before now in order to calculate vcm let our y axis to point vertically downward and x axis horizontally to the right vcm-ydirection = (1*10*cos(60) + 1*10*cos(60)) / 2 = 5 vcm-xdirection = (1*10*sin(60) - 1*10*sin(60)) / 2 = 0